OpenStudy (anonymous):

write using exponents . -3 * m* 4 *n * n A. 1mn2 B. –12mn2 C. –7mn2 D. 12mn2

OpenStudy (anonymous):

i think is c. but im not sure ??

OpenStudy (diyadiya):

Ok -3*m*4*n*n First multipply -3 * 4 What do you get?

OpenStudy (anonymous):

-12 ??

OpenStudy (diyadiya):

Right Now you've -12*m*n*n remember \[\huge x^m \times x^n= x^{m+n}\] So what will be n*n

OpenStudy (diyadiya):

\[n^1 \times n^1 =?\]

OpenStudy (anonymous):

-12mn2

OpenStudy (diyadiya):

Right \[n^1 \times n^1= n^{1+1}=n^2\] So finally you have \[-12mn^2\]

OpenStudy (anonymous):

Ohh now i got it finally :) thank you

OpenStudy (diyadiya):

You're Welcome :)